OpenStudy (abstracthaven):

i am seriously frustrated with this problem someone help me solve: find the center and radius of each circle. 2x^2+2Y^2+14x+6y+2=0

OpenStudy (mertsj):

First divide each term by 2

OpenStudy (mertsj):

\[x^2+7x+y^2+3y+1=0\]

OpenStudy (mertsj):

One-half of 7 is 7/2 and 7/2 squared is 49/4 so add 49/4 to both sides of the equation. (Also get the one over to the right side.)

OpenStudy (mertsj):

\[x^2+7x+\frac{49}{4}+y^2+3y=-1+\frac{49}{4}\]

OpenStudy (mertsj):

Now do the same thing with the 3y. One-half of 3 is 3/2 and 3/2 squared is 9/4 so add 9/4 to both sides.

OpenStudy (mertsj):

\[x^2+7x+\frac{49}{4}+y^2+3y+\frac{9}{4}=-1+\frac{49}{4}+\frac{9}{4}\]

OpenStudy (mertsj):

Now factor the left side and add the numbers on the right side together.

OpenStudy (mertsj):

\[(x+\frac{7}{2})^2+(y+\frac{3}{2})^2=\frac{27}{2}\]

OpenStudy (abstracthaven):

thanks i had everything but the right of the = sign correct I didnt divide the 2 from the original equation thanks

OpenStudy (mertsj):

yw